# QCPINN: Quantum-Classical Physics-Informed Neural Networks for Solving PDEs

### v5

@@ 4 Quantum-Classical Physics-Informed Neural Networks (QCPINN)
11:   Apply Interferometer($\psi_{1}[l]$ , $\phi_{1}[l]$)
12:   Apply Squeezing($r[l],\phi_{\text{s}}[l]$) $\triangleright$ Phase-free: $\phi_{\text{s}}[l]=0.0$
13:   Apply Interferometer($\psi_{2}[l]$ , $\phi_{2}[l]$)
14:   Apply Displacement($\alpha[l],\phi_{\text{d}}[l]$) $\triangleright$ Phase-free: $\phi_{\text{d}}[l]=0.0$
15:   Apply Nonlinearity $\Phi$ $\triangleright$ Options: Kerr, Cubic Phase, Cross-Kerr
@@
24:      Apply Beamsplitter $B(\psi[n],\phi_{\text{bs}}[n])$ $\triangleright$ Phase-free: $\phi_{\text{bs}}[n]=0.0$
@@
4.1 CV-circuit QCPINN
The CV-circuit QCPINN consists of three primary components: a preprocessor network, a CV-circuit quantum network (the QNN layer of Fig. 1), and a postprocessor network. The quantum network is adapted from [17] and is described in algorithm 1. Each layer of the quantum network executes a sequence of quantum operations.
Structuring the network as described in algorithm 1 results in a circuit depth of $(2n+3)L$, while the number of trainable quantum parameters increases to $\mathcal{O}(n^{2}L)$, where $n$ represents the number of qumodes and $L$ denotes the number of layers.
4.2 DV-circuit QCPINN
Table 2: Characteristics of the topologies used in DV-circuits in Fig. 2. $n$ is the number of qubits, and $L$ is the number of layers.
alternate $6L$ Nearest-neighbor 4(n1)L4(n-1)L (n1)L(n-1)L
cascade $(n+2)L$ Ring topology $3nL$ nLnL
cross-mesh $(n^{2}-n+4)L$ All-to-all $(n^{2}+3n)L$ $(n^{2}-n)L$
layered $6L$ Nearest-neighbor 4(n1)L4(n-1)L (n1)L(n-1)L
The DV-circuit QCPINN consists of three primary components: a preprocessor network, a DV-circuit quantum network (the QNN layer of Fig. 1), and a postprocessor network. The quantum network is a custom DV-circuit, which executes quantum operations on the preprocessed data.
Table 1 presents various configurations of our DV-circuit studied to identify a suitable embedding scheme and topology. Fig. 2 illustrates four proposed ansätze named after their topology: alternate, cascade, cross-mesh, and layered. These circuits are adapted from literature, including  [71, 72, 73, 62]. Table 2 presents features of the DV-circuits across the key descriptors, circuit depth, connectivity, number of parameters, and the number of two-qubit gates. These circuits agree with the affine transformation discussed in section 2 and adhere to the hardware-efficient ansatz (HEA) [74] defined as:
$U(\psi)\;=\;\prod_{k}U_{k}(\psi_{k})\,W_{k},$ (3)
The cross-mesh ansatz implements an all-to-all connectivity scheme, facilitating extensive interaction between qubits, which enhances its ability to explore the Hilbert space. The circuit depth scales as $\mathcal{O}(n^{2}-n+4)L$, while the number of trainable parameters is $(n^{2}+3n)L$. The use of global entanglement fosters strong correlations between qubits; however, it results in greater circuit depth, resource demands, and training difficulties [20]. This circuit’s total number of two-qubit gates is also significantly higher than in other ansätze.
The cascade ansatz balances expressivity and hardware efficiency by employing a ring topology for qubit connectivity. This design facilitates efficient entanglement while keeping the circuit depth manageable at $(n+2)L$. The circuit contains $3nL$ trainable parameters, making it more efficient than cross-mesh while maintaining reasonable expressivity. Incorporating controlled rotation gates ($CR_{X}$) improves entanglement compared to CNOT gates and increases the circuit’s capacity to explore the Hilbert space. Furthermore, the requirement for two-qubit gates is fewer than that of cross-mesh, thus lowering the overall computational cost.
The layered ansatz features a structured design that includes alternating rotation and entangling layers. Each qubit undergoes parameterized single-qubit rotations, typically utilizing $R_{X}$ and $R_{Z}$ gates. Entanglement is achieved through nearest-neighbor CNOT gates, ensuring compatibility with hardware constraints while maintaining efficient connectivity. Compared to the cross-mesh’s all-to-all connectivity or the cascade’s ring topology, the layered architecture represents a compromise between expressivity and hardware efficiency.
The alternate ansatz uses an alternating-layer structure featuring nearest-neighbor CNOT gates, ensuring effective entanglement while maintaining a manageable circuit depth of $6L$. The total number of trainable parameters in this circuit 4(n1)L4(n-1)L aligns with the layered ansatz. The entanglement structure determined by the alternating CNOT layers offers the highest level of entangling capability among the circuits studied. Consequently, the cross-mesh ansatz provides the largest parameter count and highest expressivity, whereas the alternate circuit excels in entangling capabilities.
Compared to the QNN models discussed above, we can see that the classical feed-forward neural networks (NNs) exhibit a more flexible scaling pattern, with trainable parameters $\mathcal{O}(h^{2}L)$, where $h$ denotes the number of hidden neurons (assuming a fixed width $h$ in all hidden layers) with independent selection of $h$. Therefore, while classical models depend on deep architectures and larger parameter counts, quantum networks achieve expressivity by utilizing intrinsic quantum correlations and exponential scaling of the Hilbert space.
5 Implementation and Training Details
We used PyTorch’s automatic differentiation to compute the gradients for both classical and quantum parameters. The training was carried out with a batch size of 64 and a maximum of 20,000 epochs in all case studies, except in the cavity problem, where we used 12,500 maximum iterations. The Adam optimizer was used to ensure stable convergence. These settings were uniformly applied across all study cases to provide a consistent basis for comparison. For the cavity problem, initial training data was selected using the Sobol sequence, and then random sampling with a batch size of 128 was applied to this Sobol-generated dataset during training. For all other cases, only random sampling with a batch size of 128 was used throughout the entire process.
While other training configurations, such as increasing the number of iterations or adjusting the batch size, could yield better performance, the goal is to evaluate the models under identical conditions. Given the near-infinite hyperparameter space, this controlled evaluation ensures a fair assessment of the relative strengths and limitations of each approach. It is also important to acknowledge that the quantum experiments conducted in this study were performed using emulator-based simulations rather than actual quantum hardware. These simulations inherently operate at significantly reduced speeds compared to their classical counterparts due to the substantial computational overhead associated with emulating quantum circuits on classical machines. Future implementations on actual quantum hardware would eliminate this simulation overhead, potentially revealing additional performance characteristics not captured in the current study.
In our initial implementation, we observed that the experiments on the CPU were faster than those on the GPU. This is likely due to our study’s relatively small batch size, circuit complexity, and limited number of qubits/qumodes: five qubits for DV-circuit models and two qumodes for CV-based models. Moreover, the quantum backends employed in our experiments are not optimized for GPU efficiency. Considering these limitations, we have chosen to conduct our experiments on a CPU system with 6148 CPUs (each operating at 2.40 GHz) and 192 GB of RAM. Further details on training configurations and implementation strategies are provided in the following subsections.
5.1 CV-circuit QCPINN
We used the PennyLane framework [68] with the Strawberry Fields Fock backend [75] to develop and train our CV-circuit QCPINN model. Each CV-circuit layer was configured for two qumodes with a cutoff dimension of 20. Quantum parameters were initialized using controlled random distributions, with small values for active parameters (e.g., $r$, $d$) to ensure numerical stability. Classical PINN’s layers were initialized using Xavier initialization. During training, either position quadrature (X) or number operator expectation measurements are performed on each qumode to extract continuous-valued features from the quantum state.
The CV-circuit operations are carefully managed to ensure numerical stability. For instance, squeezing operations are highly sensitive to large values, as excessive squeezing amplifies quadrature fluctuations, which increases photon number variance and truncation errors in Fock-space representations [76]. Similarly, displacement operations are constrained to prevent quantum states from exceeding computational cutoffs, as this can introduce simulation errors [77]. Likewise, the nonlinear operations such as Kerr gate can cause phase shifts that scale quadratically with photon numbers, making them highly sensitive to large values [78].
Therefore, we used a learning rate of $1.0\times 10^{-4}$ to address the model’s sensitivity to gradient updates. A ReduceLROnPlateau scheduler was implemented with a patience of 20 epochs, decreasing the learning rate by a factor of 0.5 when no improvement was observed, down to a minimum of $1.0\times 10^{-6}$. Additionally, we applied gradient clipping to prevent exploding gradients, restricting their norm to a maximum of 0.1. As shown in algorithm 1, we also scaled the training parameters by small factors to further stabilize the CV transformations.
While these settings can alleviate the instability and prevent divergence in quantum state evolution, they can also introduce the vanishing gradient problem. This issue is well-documented in classical neural network literature and occurs when excessively small initialization values diminish gradients during backpropagation, slowing or even halting the learning process [79].
5.2 DV-circuit QCPINN
We used Adam optimizer with a learning rate of 0.005 to train the DV-circuit QCPINN. A learning rate scheduler, ReduceLROnPlateau, is employed to adjust the learning rate when the loss plateaus dynamically. Specifically, the scheduler reduces the learning rate by a factor of 0.9 if no improvement is observed for 1000 consecutive epochs, promoting stability and convergence. Furthermore, the training function applies gradient clipping to prevent exploding gradients, clipping them according to their norm to ensure they do not exceed a unity value. During training, Pauli-Z expectation measurements are performed on each qubit to extract features from the quantum state.
5.3 Classical PINN Model
The classical baseline model consists of three main components: a preprocessing network, a neural network NN, and a postprocessing network (as shown in figure 1). The NN differs between two configurations: (a) In model-1, it includes two fully connected layers, each containing 50 neurons with a Tanh activation function between layers, and (b) In model-2, the NN is entirely omitted, meaning only the preprocessing and postprocessing networks are kept, simplifying the model’s structure.
These two models represent different levels of complexity, with model-1 having more layers and, consequently, a higher number of trainable parameters than model-2. We use these classical models to evaluate and compare the performance of the proposed hybrid models in figure 1. Training settings, including the optimizer and batch size, were kept consistent with the QCPINN models to allow for a fair comparison between classical and quantum-enhanced approaches. The learning rate is $0.005$.
5.4 PDEs
We implemented five distinct PDEs, which include the Helmholtz, the time-dependent 2D lid-driven cavity, the 1D wave, the Klein-Gordon, and a convection-diffusion problems. The mathematical formulation, boundary conditions, and corresponding loss function designs for each PDE are detailed in A.
@@ 5.4 PDEs
Table 3: The CV-circuit QCPINN models the performance in solving the Helmholtz and lid-driven cavity equations. There are 12 combinations in total, as in Table 1. Among these results of the model with a quadrature measurement scheme, a Kerr nonlinear function, and phase-free parameterization is reported here. The remaining 11 combinations are not mentioned as the results are more or less similar. Training convergence is shown in Fig. 3.
PDE # Parameters   Relative $\mathbf{L_{2}}$ error (%) Final loss
@@
Figure 3: The training convergence behavior of the CV-circuit QCPINN model when solving (a) Helmholtz and (b) lid-driven cavity. The terms “loss_r” and “loss_bc” refer to residual and boundary losses. The “loss_bc” shows a decreasing trend, indicating improvements in satisfying the boundary conditions; however, the overall decrease is below expectation. The “loss_r” exhibits significant fluctuations and does not display a noticeable reduction, suggesting difficulties in minimizing the residual error of the governing equation.
6 Results
We began testing our CV and DV-circuit QCPINN using two benchmark problems: Helmholtz and the lid-driven cavity.
6.1 CV-circuit QCPINN
At the beginning of our study, we implemented CV-circuit QCPINN models using 12 different combinations. This included two measurement schemes, three nonlinear functions, and two parameterization schemes (see Table 1). The results indicate that the behavior of these models is quite similar. Table 3 and Fig. 3 present the results and training convergence for the model that employs the quadrature measurement scheme, Kerr nonlinear function, and phase-free parameterization.
6.1.1 Helmholtz Equation.
For the Helmholtz Eq. 4, we used 469 trainable parameters. The results, in Table 3, show a high relative $L_{2}$ error for both the $u$ and $f$ variables, along with a significant final loss. The training convergence pattern in Fig. 3(a) highlights some challenges. For example, the boundary loss, “loss_bc”, shows a steady decrease over 10,000 iterations, suggesting that the model is gradually improving in meeting the boundary conditions. However, the residual loss, “loss_r”, remains consistently high, with substantial fluctuations throughout the training process. This indicates that the model was struggling to satisfy the differential equation itself.
6.1.2 Time-dependent 2D Lid-driven Cavity Equation.
For the more complex lid-driven cavity Eq. 5, the CV-circuit QCPINN model was implemented with 621 trainable parameters. The result, in Table 3, shows that the relative $L_{2}$ errors for both velocity and pressure variables remain high. However, the final loss value of 0.102 is significantly lower than that observed for the Helmholtz equation. Fig. 3(b) shows the training convergence, where the boundary loss exhibits some improvement, although it still experiences considerable fluctuations. The residual loss displays even greater oscillations compared to the Helmholtz case, indicating a greater challenge in accurately capturing the complex fluid dynamics governed by the Navier-Stokes equations.
@@ 6.1 CV-circuit QCPINN
6.2 DV-circuit QCPINN
The DV-circuit QCPINNs demonstrate stable performance compared to the CV-circuit across all the PDEs considered in this study. Thus, we perform a detailed analysis of the DV-circuit implementation throughout the remainder of this paper.
Our investigation began by using various DV-circuit configurations to solve two fundamental PDEs: Helmholtz and the lid-driven cavity. We systematically evaluated the performance across combinations of two distinct embedding approaches and four different circuit topologies, comparing these quantum implementations against classical models.
Table 4: Comparison of the performance of our DV-circuit QCPINN models against classical PINN approaches in solving the Helmholtz equation. The comparison metrics include the number of trainable parameters required by each model, the relative $L_{2}$ error in %, and the final training loss value achieved. $u$ and $f$ are the PDE solution and the force term.
Embedding Topology # Parameters   Relative $\mathbf{L_{2}}$ error (%) Final loss
DV-circuit amplitude alternate 781 $u$ $2.14e+01$ $55.360$
$f$ $1.21e+01$
cascade 771 $u$ $1.99e+01$ $6.372$
$f$ $4.97e+00$
cross-mesh 836 $u$ $1.77e+01$ $6.872$
$f$ $3.76e+00$
layered 781 $u$ $4.05e+01$ $39.441$
$f$ $7.54e+00$
@@ 6.2 DV-circuit QCPINN
alternate 781 $u$ $9.15e+00$ $11.456$
$f$ $4.15e+00$
cascade 771 $u$ $4.12e+00$ $3.301$
$f$ $1.64e+00$
cross-mesh 836 $u$ $1.18e+01$ $5.822$
$f$ $3.76e+00$
layered 781 $u$ $2.35e+01$ $11.512$
$f$ $5.73e+00$
Classical
@@
model-1 7851 $u$ $3.18e+01$ $20.331$
$f$ $6.40e+00$
model-2 2751 $u$ $6.72e+00$ $4.771$
$f$ $2.83e+00$
Figure 4: Training convergence history of our DV-based QCPINN models compared to best classical PINN model when solving the Helmholtz equation. The plots track the performance over 20,000 epochs, showing how QCPINN performs with two different embedding methods across four circuit topologies. For comparison, we include the classical PINN model-2, which outperforms model-1 in this case.
Figure 5: Comparison of the exact solution and the model predictions for the Helmholtz equation: the best classical PINN results (model-2), and our DV-circuit QCPINN using angle embedding (which outperforms amplitude embedding) with cascade topology (which outperforms the other three topologies). The first row shows the PDE solution $u$, while the second row shows the force field ($f$).
6.2.1 Helmholtz Equation.
Table 4 presents the performance of the DV-circuit QCPINN models in solving the Helmholtz Eq. 4 and comparing the results with classical models. The results show that the cascade topology consistently outperformed other topologies, achieving the lowest final training loss and low relative $L_{2}$ testing errors. In contrast, the layered and alternate circuit topologies show higher final training losses, especially the layered topology with amplitude embedding.
Fig. 4 compares the training loss convergence of DV-circuit QCPINN models to the classical PINN. The results indicate that DV-based quantum models display controlled convergence with relatively stable training dynamics. Moreover, the angle embedded cascade topology demonstrates the most competitive performance, achieving a loss level close to the best classical model while utilizing approximately 72% fewer parameters.
Fig. 5 visualizes the exact and predicted solutions for the Helmholtz equation. It also shows the prediction errors of both the best classical model and the angle-cascade QCPINN. While the classical model captures the overall structure, it exhibits noticeable discrepancies, as shown in the absolute error maps, which reveal significant deviations near the boundaries and at regions of sharp local gradients. In contrast, the predictions from the angle-cascade QCPINN model closely align with the exact solutions while using fewer parameters than the classical approach.
Table 5: Comparison of the performance of our DV-circuit QCPINN models against classical PINN approaches in solving the lid-driven cavity equation. The comparison metrics include the number of trainable parameters required by each model, the relative $L_{2}$ error in %, and the final training loss value achieved. $u$, $v$, and $p$ are the velocity components and the pressure values, respectively.
Embedding Topology # Parameters   Relative $\mathbf{L_{2}}$ error (%) Final loss
DV
amplitude
alternate 933 $u$ $6.441e+01$ 0.127
$v$ $8.439e+01$
pp $6.102e+01$
cascade 928 $u$ $6.670e+01$ 0.123
$v$ $8.409e+01$
pp $6.365e+01$
cross-mesh 988 $u$ $7.073e+01$ 0.109
$v$ $4.893e+01$
pp $7.096e+01$
layered 933 $u$ $5.690e+01$ 0.121
$v$ $9.476e+01$
pp $7.469e+01$
angle
alternate 933 $u$ $4.041e+01$ 0.109
$v$ $4.683e+01$
pp $4.746e+01$
cascade 928 $u$ $2.745e+01$ 0.083
$v$ $2.081e+01$
pp $4.791e+01$
cross-mesh 988 $u$ $2.035e+01$ 0.084
v $2.071e+01$
p $4.011e+01$
layered 933 $u$ $4.572e+01$ 0.129
v $6.628e+01$
p $7.885e+01$
Classical
PINN
model-1 8003 $u$ $2.103e+01$ 0.087
$v$ $3.027e+01$
pp $2.317e+01$
model-2 2903 $u$ $3.830e+01$ 0.073
$v$ $2.924e+01$
pp $4.106e+01$
Figure 6: Training convergence history of our DV-based QCPINN models compared to the best classical PINN model (model-2) when solving the lid-driven cavity equation. The plots track the performance over 12500 epochs, showing how QCPINN performs with two different embedding methods across four circuit topologies.
Figure 7: Comparison of the reference FEM solution and the model predictions for the lid-driven cavity equation: the classical PINN model-2 (which outperforms model-1), and our DV-circuit QCPINN using angle embedding (which outperforms amplitude embedding) with cascade topology (which outperforms the other three topologies). The first and second rows show the velocity fields ($u$ and $v$), while the third row shows the pressure field ($p$).
6.2.2 Time-dependent 2D Lid-driven Cavity Equation.
Table 5 presents the performance of the QCPINN methods for solving the lid-driven cavity Eq. 5. Among the DV-circuit QCPINN models, cascade achieves the best overall performance, demonstrating good convergence properties with the lowest final training loss. The models with angle embedding and cross-mesh topology also perform well, achieving the lowest final training loss and highlighting their effectiveness in this setting. Specifically, the angle-cascade quantum model achieves comparable performance to the best classical model while requiring only 12% of the parameters.
Fig. 6 compares the training loss convergence of various DV-circuit QCPINN models with a classical PINN while solving the lid-driven cavity equation. The loss curves indicate that all DV-based models achieve stable convergence with minimal oscillations. Similar to Fig. 6, the angle embedding with the cascade topology circuit in Fig. 4 closely matches the loss level of the best classical model.
Fig. 7 visualizes the results obtained using the finite element method solution alongside the predicted solutions for the lid-driven cavity Eq. 5. It also shows the prediction errors for both the best classical model and the angle-cascade QCPINN. The classical PINN model and the angle-cascade QCPINN effectively capture the main flow features. The absolute error distributions (third and fifth rows) reveal distinct patterns in both velocity components and pressure fields. Both models exhibit higher errors near the top moving lid for velocity components, where the flow experiences the most significant shear.
6.2.3 Generalizing Our Solution.
Table 6: Comparison of our DV-circuit QCPINN using angle embedding (which outperforms amplitude embedding) with cascade topology (which outperforms the other three topologies) with the two classical PINN models for solving wave, Klein-Gordon, and convection-diffusion equations. The comparison metrics include the number of trainable parameters required by each model, the relative $L_{2}$ error in %, and the final training loss value achieved. $u$ and $f$ are the reference solutions and the source term, respectively.
PDEs Method
# Parameters
Relative $\mathbf{L_{2}}$ error (%)
Final loss
wave Eq. 6 angle-cascade 771 $u$ $1.02e+01$ $0.083$
Classical
PINN
model-1 $7851$ $u$ $2.06e+01$ $0.128$
model-2 2751 $u$ $8.68e+00$ $0.027$
@@
angle-cascade 771 $u$ $1.17e+01$ $2.727$
$f$ $3.17e+00$
Classical
@@
angle-cascade 821 $u$ $1.20e+01$ $0.008$
$f$ $4.38e+00$
Classical

### v6

@@ B.1 Algorithm
11:   Interferometer($\psi_{1}[l]$ , $\phi_{1}[l]$)
12:   Squeezing($r[l],\phi_{\text{s}}[l]$) $\triangleright$ Phase-free: $\phi_{\text{s}}[l]=0.0$
13:   Interferometer($\psi_{2}[l]$ , $\phi_{2}[l]$)
14:   Displacement($\alpha[l],\phi_{\text{d}}[l]$) $\triangleright$ Phase-free: $\phi_{\text{d}}[l]=0.0$
15:   Nonlinearity $\Phi$ $\triangleright$ Options: Kerr, Cubic Phase, Cross-Kerr
@@
24:      Beamsplitter $B(\psi[n],\phi_{\text{bs}}[n])$ $\triangleright$ Phase-free: $\phi_{\text{bs}}[n]=0.0$
@@
B.2 Results
@@
Table 4: The performance of CV-Circuit QCPINN models in solving the Helmholtz and lid-driven Cavity equations. There are 12 combinations reported in Table 1, however, results with the quadrature measurement scheme, Kerr nonlinear function, and phase-free parameterization are reported here. The remaining 11 combinations are not mentioned as the results are more or less similar. $N_{p}$: number of parameters, $L_{2}$: relative $L_{2}$ error in %, $\mathcal{L}_{\text{final}}$: final loss. Training convergence is shown in Fig. 9.
PDE $N_{p}$ $L_{2}$ $\mathcal{L}_{\text{final}}$
@@
Figure 9: The training convergence of the CV-Circuit QCPINN model when solving Helmholtz and lid-driven Cavity. The terms “loss_r” and “loss_bc” refer to residual and boundary losses. The “loss_bc” shows a decreasing trend, indicating improvements in satisfying the boundary conditions; however, the overall decrease is below expectation. The “loss_r” exhibits significant fluctuations and does not display a noticeable reduction, suggesting difficulties in minimizing the residual error of the governing equation.
Helmholtz Equation. For the Helmholtz Eq. 5, we used 469 trainable parameters. The results, in Table 4, show a high relative $L_{2}$ error for both the $u$ and $f$ variables, along with a significant final loss. The training convergence pattern in Fig. 9(a) highlights some challenges. For example, the boundary loss, “loss_bc”, shows a steady decrease over 10,000 iterations, suggesting that the model is gradually improving in meeting the boundary conditions. However, the residual loss, “loss_r”, remains consistently high, with substantial fluctuations throughout the training process. This indicates that the model was struggling to satisfy the differential equation itself.
Time-dependent 2D Lid-driven Cavity Equation. For the more complex lid-driven Cavity Eq. 6, the CV-Circuit QCPINN model was implemented with 621 trainable parameters. The result, in Table 4, shows that the relative $L_{2}$ errors for both velocity and pressure variables remain high. However, the final loss value of 0.102 is significantly lower than that observed for the Helmholtz equation. Fig. 9(b) shows the training convergence, where the boundary loss exhibits some improvement, although it still experiences considerable fluctuations. The residual loss displays even greater oscillations compared to the Helmholtz case, indicating a greater challenge in accurately capturing the complex fluid dynamics governed by the Navier-Stokes equations.
@@
Appendix C DV-Circuit QCPINN
Table 5: Results from all configurations of QCPINN (DV-Circuit) models and PINN models for solving the Helmholtz equation. $N_{p}$: number of trainable parameters, $L_{2}$: relative error (in %), $L_{\text{final}}$: final loss, T: time per iteration (in sec), M: peak memory consumption (in MB), $u$: reference solution, and $f$ source term.
Embedding Topology $N_{p}$ $L_{2}$ $\mathcal{L}_{\text{final}}$ T M
QCPINN
(DV-Circuit)
Amplitude Alternate 772 $u$ $31.98\pm 9.51$ $24.42\pm 10.91$ $0.26\pm 0.00$ $420.0\pm 0.0$
$f$ $7.70\pm 1.42$
Cascade 771 $u$ $21.33\pm 8.19$ $35.49\pm 29.26$ $0.28\pm 0.00$ $420.0\pm 0.0$
$f$ $8.03\pm 2.83$
Cross-mesh 796 $u$ $29.39\pm 17.84$ $23.29\pm 16.08$ $0.31\pm 0.0$ $440.0\pm 0.0$
$f$ $7.30\pm 3.21$
Layered 776 $u$ $20.95\pm 7.46$ $18.16\pm 9.18$ $0.29\pm 0.03$ $420.0\pm 0.0$
$f$ $4.96\pm 0.92$
@@ Appendix C DV-Circuit QCPINN
Alternate 772 $u$ $10.74\pm 5.54$ $4.87\pm 3.19$ $0.18\pm 0.0$ $420.0\pm 0.0$
$f$ $3.42\pm 0.8$
Cascade 771 $u$ $\mathbf{6.69\pm 1.51}$ $4.52\pm 1.86$ $0.17\pm 0.00$ $420.0\pm 0.0$
$f$ $\mathbf{2.86\pm 0.54}$
Cross-mesh 796 $u$ $7.24\pm 1.45$ $3.69\pm 1.30$ $0.25\pm 0.02$ $430.0\pm 0.0$
$f$ $3.09\pm 0.70$
Layered 776 $u$ $8.19\pm 10.24 11$$1.63\pm 8.73$ $0.19\pm 0.00$ $420.0\pm 0.0$
$f$ $4.88\pm 1.42$
PINN Model-1 7851 $u$ $21.24\pm 12.13$ $16.62\pm 13.85$ $0.0041\pm 0.0$ $340.0\pm 0.0$
$f$ $5.59\pm 2.56$
Model-2 2751 $u$ $12.12\pm 7.17$ $5.04\pm 3.12$ $0.0027\pm 0.0$ $340.0\pm 0.0$
$f$ $3.23\pm 1.19$
Figure 10: Training convergence history of all QCPINN (DV-Circuit) models compared to the best classical PINN model when solving the Helmholtz equation.
Table 6: Results from all configurations of QCPINN (DV-Circuit) models and PINN models for solving the Cavity problem. $N_{p}$: number of trainable parameters, $L_{2}$: relative error (in %), $L_{\text{final}}$: final loss, T: time per iteration (in sec), M: peak memory consumption (in MB), $u$ and $v$: velocity, and $p$: pressure.
Embedding Topology $N_{p}$ $L_{2}$ $\mathcal{L}_{\text{final}}$ T M
QCPINN
Amplitude
Alternate 924 $u$ $54.40\pm 7.56$ $0.08\pm 0.01$ $0.57\pm 0.0$ $570.0\pm 0.0$
$v$ $83.07\pm 1.04$
$p$ $55.53\pm 4.38$
Cascade 928 $u$ $59.70\pm 2.61$ $0.10\pm 0.03$ $0.69\pm 0.07$ $580.0\pm 0.0$
$v$ $80.15\pm 3.38$
$p$ $56.44\pm 7.54$
Cross-mesh 948 $u$ $39.95\pm 0.00$ $0.07\pm 0.00$ $0.65\pm 0.0$ $620.0\pm 0.01$
$v$ $49.62\pm 1.10$
$p$ $57.13\pm 2.00$
Layered 928 $u$ $59.92\pm 3.08$ $0.09\pm 0.03$ $0.60\pm 0.02$ $590.0\pm 0.0$
$v$ $82.67\pm 3.33$
$p$ $51.92\pm 1.99$
Angle
Alternate 924 $u$ $26.13\pm 2.00$ $0.05\pm 0.00$ $0.37\pm 0.0$ $540.0\pm 0.0$
$v$ $43.86\pm 9.00$
$p$ $43.86\pm 4.10$
Cascade 923 $u$ $25.73\pm 2.49$ $0.10\pm 0.02$ $0.36\pm 0.0$ $540.0\pm 0.0$
$v$ $38.00\pm 3.76$
$p$ $42.65\pm 8.69$
Cross-mesh 948 $u$ $27.30\pm 4.87$ $0.08\pm 0.03$ $0.47\pm 0.0$ $590.0\pm 0.0$
v $39.88\pm 4.58$
p $44.99\pm 9.24$
Layered 928 $u$ $59.87\pm 11.71$ $0.1\pm 0.17$ $4.01\pm 0.02$ $560.0\pm 0.0$
v $85.41\pm 8.22$
p $73.36\pm 13.45$
PINN Model-1 8003 $u$ $\mathbf{21.61\pm 6.68}$ $06\pm 0.02 0.0$$01\pm 0.0$ $440.0\pm 0.0$
$v$ $\mathbf{34.24\pm 2.76}$
$p$ $\mathbf{34.82\pm 5.25}$
Model-2 2903 $u$ $23.03\pm 2.45$ $06\pm 0.02 0.0$$01\pm 0.0$ $435\pm 0.0$
$v$ $38.81\pm 2.73$
$p$ $71.51\pm 4.79$
Figure 11: Training convergence history of all QCPINN (DV-Circuit) models compared to the best classical PINN model when solving the Cavity problem.
Table 7: Results from selected configurations of QCPINN (DV-Circuit) models and two PINN models. $N_{p}$: number of trainable parameters, $L_{2}$: relative error (in %), $L_{\text{final}}$: final loss, T: time per iteration (in sec), M: peak memory consumption (in MB), $u$: velocity, and $f$: source terms.
PDEs Method $N_{p}$ $L_{2}$ $\mathcal{L}_{\text{final}}$ T M
Wave Eq. 7 QCPINN Angle-Alternate 772 $u$ $39.96\pm 7.85$ $0.78\pm 0.46$ $0.15\pm 0.0$ $405.0\pm 0.0$
Angle-Layered 776 $u$ $44.76\pm 0.28$ $1.13\pm 0.26$ $0.17\pm 0.0$ $419.0\pm 0.0$
Angle-Cascade 771 $u$ $16.74\pm 5.41$ $0.10\pm 0.03$ $0.25\pm 0.06$ $446.0\pm 0.0$
Angle-Cross-mesh 796 $u$ $12.44\pm 4.42$ $0.09\pm 0.03$ $0.38\pm 0.08$ $469.0\pm 0.0$
@@
Model-1 $7851$ $u$ $20.94\pm 1.70$ $15\pm 0.04 0.0$$02\pm 0.03$ $353.0\pm 0.0$
Model-2 2751 $u$ $\mathbf{11.13\pm 1.78}$ $05\pm 0.01 0.0$$02\pm 0.02$ $351\pm 0.00$
@@
QCPINN Angle-Alternate 772 $u$ $15.54\pm 5.56$ $0.51\pm 0.37$ $0.16\pm 0.0$ $406.0\pm 0.0$
$f$ $1.99\pm 0.96$
Angle-Layered 776 $u$ $14.86\pm 4.78$ $0.51\pm 0.41$ $0.17\pm 0.0$ $409.0\pm 0.0$
$f$ $2.33\pm 1.10$
Angle-Cascade 771 $u$ $\mathbf{4.24\pm 1.88}$ $0.27\pm 0.10$ $0.25\pm 0.03$ $447.0\pm 0.0$
$f$ $\mathbf{1.72\pm 0.50}$
Angle-Cross-mesh 796 $u$ $5.64\pm 1.87$ $0.34\pm 0.21$ $0.42\pm 0.11$ $468.0\pm 0.0$
$f$ $1.65\pm 0.38$
@@
QCPINN Angle-Alternate 822 $u$ $5.06\pm 2.08$ $0.001\pm 0.001$ $0.13\pm 0.0$ $403.0\pm 0.0$
$f$ $3.22\pm 0.65$
Angle-Layered 826 $u$ $5.10\pm 3.22$ $0.01\pm 0.01$ $0.15\pm 0.0$ $410.0\pm 0.0$
$f$ $3.40\pm 1.62$
Angle-Cascade 821 $u$ $5.04\pm 1.43$ $0.001\pm 0.001$ $0.27\pm 0.06$ $445.0\pm 0.0$
$f$ $3.03\pm 1.00$
Angle-Cross-mesh 846 $u$ $\mathbf{4.54\pm 2.17}$ $0.001\pm 0.001$ $0.42\pm 0.11$ $468.0\pm 0.0$
$f$ $\mathbf{2.63\pm 1.61}$